# Privacy-preserving Blockchain-enabled Parametric Insurance via Remote Sensing and IoT

### v2

@@ VI-A Remote Sensing Model for Bushfire Detection
$(\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i})=\mathbf{n}^{-}_{i}(\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i})+\mathbf{\theta}^{-}_{i},&\mbox{for\ }1\leq i\leq n\\ (\mathbf{r}^{+}_{i}-\mathbf{s}^{+}_{i})=\mathbf{n}^{+}_{i}(\mathbf{r}^{+}_{i}+\mathbf{s}^{+}_{i})+\mathbf{\theta}^{+}_{i},&\mbox{for\ }1\leq i\leq n\\ \theta_{\max}-\sum^{n}_{i=1}(\mathbf{\theta}^{-}_{i})^{2}-\sum^{n}_{i=1}(\mathbf{\theta}^{+}_{i})^{2}=\theta_{d}>0\\ G=\sum^{n}_{i=1}{\mathds{1}}(\mathbf{n}^{-}_{i}-\mathbf{n}^{+}_{i}\geq\kappa)-\epsilon\end{array}\right.$
@@
$\mathbf{i}_{i}(1-\mathbf{i}_{i})=0,&\mbox{for\ }1\leq i\leq n\\ \mathbf{e}_{i,j}(\mathbf{e}_{i,j}-2^{j-1})=0,&\mbox{for\ }1\leq i\leq n,1\leq j\leq k\\ \sum_{j=1}^{k}\mathbf{e}_{i,j}-\mathbf{i}_{i}(\mathbf{n}^{-}_{i}-\mathbf{n}^{+}_{i}-\kappa)=0,&\mbox{for\ }1\leq i\leq n\\ G+\epsilon\phantom{1}=\sum^{n}_{i=1}\mathbf{i}_{i}&\end{array}\right.$
@@
Evidently, the above constraint system can be proved by Sonic zk-SNARK protocol [11]. We specify the settings of Sonic-specific vectors $({\bf a},{\bf b},{\bf c})$ and $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ with a more detailed explanation in Appendix VIII-E. There are in total $27n+(n+2)k+$ 2 linear constraints, and $10n+(n+2)k$ multiplication constraints, where $n$ is the image size and $k$ is the length of bit indices.
@@ VII Implementation and Evaluation
    Enhanced Verifier: The verifier of enhanced protocol (${\tt S}_{\tt EV}$) in Table XIII with batch verification.
@@ VII-A Performance of Protocol
    Num. Linear Constraints: The number of linear constraints included in the constraint system. The settings of vectors ($\mathbf{k}_{q}$, $\mathbf{u}_{q}$, $\mathbf{v}_{q}$ and $\mathbf{w}_{q}$) in the linear constraints are provided in Table XIV.
@@
    Num. Multiplication Constraints: The number of multiplication constraints in the constraint system. The settings of vectors ($\mathbf{a}$, $\mathbf{b}$ and $\mathbf{c}$) in the multiplication constraints are provided in Table XV.
@@
[…]
    Evaluation based on the implementation of the protocol in Table XIII.
@@
Global Contract Deployment Cost Gas 2667K 2677K 2334K
@@
Individual Contract Deployment Cost Gas 156K 156K 156K
@@ VIII Conclusion
The practical implications of our framework for insurers and insurees are significant, as it streamlines insurance processes, reduces operational costs, and fosters trust between all parties involved. Moreover, the framework’s potential extension to other parametric insurance applications, such as health, flood, and crop insurance, underscores its versatility and adaptability. Our research contributes to the ongoing efforts to improve the efficiency, transparency, and security of insurance practices and provides a foundation for further research in this field.
@@ VIII-A Restricted KZG Polynomial Commitment
We present a restricted version of KZG polynomial commitment scheme (denoted by ${\tt rKZG}$) in Table XI that precludes a committed polynomial with a non-zero constant term. Given a Laurent polynomial $f[X]=\sum_{i=-d}^{d}a_{i}X^{i}$, this restricted polynomial commitment scheme does not allow the input $a_{0}$, and hence, a committed polynomial must have a zero constant term. Note that our scheme simplifies the one in [11], which also considers the degree of $f[X]$ bounded by a known constant less than $d$.
@@ VIII-B Restricted Polynomial Commitment with Batch Verification
  • $\textsf{\footnotesize Setup}_{\tt rKZGb}$: We suppose that there is a trusted party, who takes the security parameter ${\lambda}$ and generates $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{\mathcal{S}}$ with bilinear pairing ${\tt e}$. Then, it randomly selects $g\in{\mathbb{G}_{1}},h\in{\mathbb{G}_{2}},({\alpha,x})\xleftarrow{\$}{\mathbb{F}_{p}}\backslash\{0,1\}$. Next, set the structured reference string as:
@@
$\gamma_{i}[X]\triangleq\sum_{z\in{\mathcal{S}}_{i}}f_{i}[z]\cdot\Big{(}\frac{\prod_{z^{\prime}\in{\mathcal{S}}_{i}\backslash\{z\}}(X-z^{\prime})}{\prod_{z^{\prime}\in{\mathcal{S}}_{i}\backslash\{z\}}(z-z^{\prime})}\Big{)}$
@@
We describe the restricted KZG polynomial commitment scheme with batch verification (${\tt rKZGb}$) in Table XII. Note that $\textsf{\footnotesize BatchVerify}_{\tt rKZGb}$ is an interactive process, which can be converted to a non-interactive one by Fiat-Shamir heuristic.
VIII-C Proof of Knowledge Soundness for ${\tt rKZGb}$
@@
Given ${\tt srs}\leftarrow\textsf{\footnotesize\em Setup}_{\tt rKZGb}(\lambda,{\alpha,x})$, where $({\alpha,x})\xleftarrow{\$}{\mathbb{F}_{p}}\backslash\{0,1\}$, the probability that any algebraic adversary $\mathcal{A}$ can output $({\alpha,x})$ efficiently from ${\tt srs}$ is negligible $\mathbf{\epsilon}(\lambda)$.
@@
Given polynomial $F[X]=\sum_{i=1}^{K}\beta^{i-1}\cdot f_{i}[X]$, where $f_{i}[X]$ is a polynomial over a finite field $\mathbb{F}$ with a degree bounded in $[-d,d]$, and $Z[X]$ that decomposes to distinct linear factors over $\mathbb{F}$. Suppose for some $j\in\{1,...,K\}$, $f_{j}[X]$ is indivisible by $Z[X]$, then with probability at least $1-\frac{K}{|\mathbb{F}|}$, $F[X]$ is also indivisible by $Z[X]$.
By Lemma 3, we obtain that $\hat{f}[X]$ is indivisible by $Z_{{\mathcal{S}}}[X]$ with probability at least $1-\frac{K}{|\mathbb{F}_{p}|}$. If $\hat{f}[X]$ is divisible by $Z_{{\mathcal{S}}}[X]$, then $\mathcal{A}$ wins, but with a negligible probability $\frac{K}{|\mathbb{F}_{p}|}$, as $|\mathbb{F}_{p}|$ increases with $\lambda$. Suppose $\hat{f}[X]$ is indivisible by $Z_{{\mathcal{S}}}[X]$ which we call Assumption ($\ast$). Then $\mathcal{A}$ provides $\tilde{p}[X]$ in $\pi_{1}$, but it is certain that $\hat{f}[X]\neq\tilde{p}[X]\cdot Z_{{\mathcal{S}}}[X]$.
@@
We define a “real pairing check” with a group element $F$ that means checking $F$ by invoking some pairing equation with ${\tt e}\langle\cdot,\cdot\rangle$, whereas the corresponding “ideal pairing check” means checking $F$ instead with the vector $(a_{t})_{t=1}^{N}$ where $F=\prod_{t=1}^{N}{\tt srs}_{t}^{a_{t}}$ by some linear equation. We adopt the following Lemma from [19].
@@
$\alpha\cdot\tilde{w}_{\mu}[x]\cdot(x-\mu)$ =?${=}\alpha\cdot\hat{f}[\mu]-\alpha\cdot\tilde{p}[x]\cdot Z_{\mathcal{S}}[x]$ (10)
$\Rightarrow$ $\tilde{w}_{\mu}[x]\cdot(x-\mu)$ =?${=}\hat{f}[\mu]-\tilde{p}[x]\cdot Z_{\mathcal{S}}[x]$ (11)
@@
TABLE XIII: Enhanced Interactive zk-SNARK Protocol with Heterogeneous Data Sources and Batch Verification (${\tt S}_{\tt EV}$)
@@
    Prover $\Rightarrow$ Verifier: //${\mathcal{S}}_{X},\hat{k},\hat{s},\hat{s}_{1},\hat{s}_{2}$ computing is outsourced to prover:
@@
    $(\tilde{R},\gamma_{\tilde{R}})$ $\leftarrow{\textsf{\scriptsize Commit}_{\tt rKZGb}}({\tt srs},\tilde{r}[X,1])$
@@
TABLE XIV: Settings of vectors $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ in the constraint system for parametric bushfire insurance. $\mathbf{I}(i,j)$ represents a vector of length $(10n+(n+2)k)$ for which every entry are 00 except $i_{th},j_{th}$ entries which are $1$.
@@
$\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i},i\in\{1,...,n\}$ 00 $\mathbf{I}(6n+(n+2)k+i,7n+(n+2)k+i)$ $\mathbf{-I}(i)$ $\mathbf{I}$
$\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i}-\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$ 00 $\mathbf{I}(6n+(n+2)k+i)-I(7n+(n+2)k+i,3n+i)$ $\mathbf{I}$ $\mathbf{-I}(i)$
$\mathbf{r}^{+}_{i}+\mathbf{s}^{+}_{i},i\in\{1,...,n\}$ 00 $\mathbf{I}(8n+(n+2)k+i,9n+(n+2)k+i)$ $\mathbf{-I}(n+i)$ $\mathbf{I}$
$\mathbf{r}^{+}_{i}-\mathbf{s}^{+}_{i}-\mathbf{\theta}^{+}_{i},i\in\{1,...,n\}$ 00 $\mathbf{I}(8n+(n+2)k+i)-I(9n+(n+2)k+i,4n+i)$ $\mathbf{I}$ $\mathbf{-I}(n+i)$
@@
$\sum^{k}_{j=1}e_{i,j},i\in\{1,...,n\}$ 00 $\mathbf{I}(6n+(i-1)k+1,...,6n+ik)$ $\mathbf{I}$ $\mathbf{-I}(2n+i)$
@@
Checking $\mathbf{i}_{i}$ in $\mathbf{a},\mathbf{b},i\in\{1,...,n\}$ $1$ $\mathbf{I}(5n+i)$ $\mathbf{I}(5n+i)$ $\mathbf{I}$
@@
VIII-D Enhanced Sonic Protocol with Batch Verification
By replacing the original restricted KZG polynomial commitment scheme by the one with batch verification, we present an enhanced version of zk-SNARK protocol (${\tt S}_{\tt EV}$) in Table XIII. To convert the interactive protocol ${\tt S}_{\tt EV}$ to be non-interactive, we can employ Fiat-Shamir heuristic to replace the verifier-supplied random challenges $(y,\beta,z)$ by hash values from the previous commitments.
VIII-E Implementation of Bushfire Insurance by Sonic Protocol
In this section, we describe how to encode the bushfire detection model in Sec.VI-A into a Sonic-compatible constraint system. We specify the detailed settings of Sonic-specific vectors $({\bf a},{\bf b},{\bf c})$ and $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ in Tables XIV and XV. We assume image size $n$ and $k$ is the length of bit indices.
We first construct the input vectors $({\bf a},{\bf b},{\bf c})$. Row $1$ to $n$ of these vectors are $\mathbf{n}^{-}$, $(\mathbf{r}^{-}+\mathbf{s}^{-})$ and $(\mathbf{r}^{-}-\mathbf{s}^{-}-\mathbf{\theta}^{-})$ respectively. Because ${\bf a}\cdot{\bf b}={\bf c}$, it verifies $(\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i})=\mathbf{n}^{-}_{i}(\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i})+\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$. Similarly, we construct row $n+1$ to $6n$ to verify the rest of the multiplication gates, which are shown in Table XV. Row $15n+$ 1 to row $6n$ verifies $\bf i$ is binary and row $16n+$ 1 to $6n+(n+2)k$ verifies $\bf e,\theta_{d},G$ are a non-negative. Finally, row $16n+(n+2)k+$ 1 to the end is input from the data source.
Next we construct $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$. The first step is to verify entries in $({\bf a},{\bf b},{\bf c})$ are consistent with the input from data sources. For example, row $1$ t $n$ of $\bf b$ are $\mathbf{r}^{-}+\mathbf{s}^{-}$, while $\mathbf{r}^{-}$ and $\mathbf{s}^{-}$ are row $6n+(n+2)k+1,...,8n+(n+2)k$ of $\bf a$. Then the first constraint in Table XIV checks $\bf a_{6n+(n+2)k+i}+\bf a_{7n+(n+2)k+i}=\bf b_{i},i\in{1,...,n}$. The next step is to encode the sum gates in the circuits. For example, checking $\sum^{n}_{i=1}\mathbf{i}_{i}-G=\epsilon$, where $\bf i$ locates in row $15n+$ 1 to $6n$, and $G$ in row $16n+(n+1)k+$ 1 to $6n+(n+2)k$ in binary decomposition form, and $\epsilon$ in $k_{q}$ as a public input. The last part checks binary decompositions are correct.

### v3

@@ VI-A Remote Sensing Model for Bushfire Detection
[…]
@@
[…]
@@
Evidently, the above constraint system can be proved by Sonic zk-SNARK protocol [8]. We specify the settings of Sonic-specific vectors $({\bf a},{\bf b},{\bf c})$ and $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ with a more detailed explanation in [29]. There are in total $7n+(n+2)k+2$ linear constraints, and $10n+(n+2)k$ multiplication constraints, where $n$ is the image size and $k$ is the length of bit indices.
@@ VII Implementation and Evaluation
    Enhanced Verifier: The verifier of enhanced protocol (${\tt S}_{\tt EV}$) with batch verification (see [29]).
@@ VII-A Performance of Protocol
    Num. Linear Constraints: The number of linear constraints included in the constraint system.
@@
    Num. Multiplication Constraints: The number of multiplication constraints in the constraint system.
@@
  • $\dagger$}
    Evaluation based on the implementation of the protocol in Table VI using the commitment scheme in Table XIV.
@@
Global Contract
Deployment
Cost
Gas 2667K 2677K 2334K
@@
Individual Contract
Deployment
Cost
Gas 156K 156K 156K
@@ VIII-A Restricted KZG Polynomial Commitment
We present a restricted version of KZG polynomial commitment scheme (denoted by $\tt rKZG$}) in Table XIII that precludes a committed polynomial with a non-zero constant term. Given a Laurent polynomial $f[X]=\sum_{i=-d}^{d}a_{i}X^{i}$, this restricted polynomial commitment scheme does not allow the input $a_{0}$, and hence, a committed polynomial must have a zero constant term. Note that our scheme simplifies the one in [8], which also considers the degree of $f[X]$ bounded by a known constant less than $d$.
@@ VIII-B Restricted Polynomial Commitment with Batch Verification
[…]
@@
[…]
@@
We describe the restricted KZG polynomial commitment scheme with batch verification ($\tt rKZGb$}) in Table XIV. Note that $\textsf{\footnotesize BatchVerify}_{\tt rKZGb}$ is an interactive process, which can be converted to a non-interactive one by Fiat-Shamir heuristic.
Appendix
VIII-C Restricted KZG Polynomial Commitment
We present a restricted version of KZG polynomial commitment scheme (denoted by $\tt rKZG$}) in Table XIII that precludes a committed polynomial with a non-zero constant term. Given a Laurent polynomial $f[X]=\sum_{i=-d}^{d}a_{i}X^{i}$, this restricted polynomial commitment scheme does not allow the input $a_{0}$, and hence, a committed polynomial must have a zero constant term. Note that our scheme simplifies the one in [8], which also considers the degree of $f[X]$ bounded by a known constant less than $d$.
TABLE XIII: Restricted KZG Polynomial Commitment Scheme ($\tt rKZG$})
{longfbox}
    Note that ${\tt g}^{\alpha}$ is explicitly removed from $\tt srs$}.
    $\pi\leftarrow{\tt g}^{q[x]}=\prod^{d}_{i=-d}({\tt g}^{{x}^{i}})^{b_{i}}$
VIII-D Restricted Polynomial Commitment with Batch Verification
TABLE XIV: Restricted KZG Polynomial Commitment Scheme
with Batch Verification ($\tt rKZGb$})
{longfbox}
[…]
    $\pi_{1}\leftarrow{\tt g}^{p[x]},\ \pi_{2}\leftarrow{\tt g}^{w_{\mu}[x]}$
    1. 1.
    2. 2.
      The verifier receives $\pi_{1}$ from the prover.
    3. 3.
    4. 4.
      The verifier receives $\pi_{2}$ from the prover.
    5. 5.
      The verifier checks the following pairing equation:
[…]
Note that
$\ell_{\mu}[X]$ $\triangleq\hat{f}_{\mu}[X]-\hat{f}[X]$
Next, we derive the following equation:
We describe the restricted KZG polynomial commitment scheme with batch verification ($\tt rKZGb$}) in Table XIV. Note that $\textsf{\footnotesize BatchVerify}_{\tt rKZGb}$ is an interactive process, which can be converted to a non-interactive one by Fiat-Shamir heuristic.
VIII-E Proof of Knowledge Soundness for $\tt rKZGb$}
@@
[…]
@@
Given polynomial $F[X]=\sum_{i=1}^{K}\beta^{i-1}\cdot f_{i}[X]$, where $f_{i}[X]$ is a polynomial over a finite field $\mathbb{F}$ with a degree bounded in [d,d][-d,d], and $Z[X]$ that decomposes to distinct linear factors over $\mathbb{F}$. Suppose for some $j\in\{1,...,K\}$, $f_{j}[X]$ is indivisible by $Z[X]$, then with probability at least $1-\frac{K}{|\mathbb{F}|}$, $F[X]$ is also indivisible by $Z[X]$.
By Lemma 3, we obtain that $\hat{f}[X]$ is indivisible by $Z_{{\mathcal{S}}}[X]$ with probability at least $1-\frac{K}{|\mathbb{F}_{p}|}$. If $\hat{f}[X]$ is divisible by $Z_{{\mathcal{S}}}[X]$, then $\mathcal{A}$ wins, but with a negligible probability $\frac{K}{|\mathbb{F}_{p}|}$, as $|\mathbb{F}_{p}|$ increases with $\lambda$. Suppose $\hat{f}[X]$ is indivisible by $Z_{{\mathcal{S}}}[X]$ - which we call Assumption ($\ast$). Then $\mathcal{A}$ provides $\tilde{p}[X]$ in $\pi_{1}$, but it is certain that $\hat{f}[X]\neq\tilde{p}[X]\cdot Z_{{\mathcal{S}}}[X]$.
@@
We define a “real pairing check” with a group element $F$ that means checking $F$ by invoking some pairing equation with $\tt e$$\langle\cdot$$\cdot\rangle,$ whereas the corresponding “ideal pairing check” means checking $F$ instead with the vector $(a_{t})_{t=1}^{N}$ where $F=\prod_{t=1}^{N}{\tt srs}_{t}^{a_{t}}$ by some linear equation. We adopt the following Lemma from [16].
@@
$\alpha\cdot\tilde{w}_{\mu}[x]\cdot(x-\mu)$ $\overset{?}{=}\alpha\cdot\hat{f}[\mu]-\alpha\cdot\tilde{p}[x]\cdot Z_{\mathcal{S}}[x]$ (15)
$\Rightarrow$ $\tilde{w}_{\mu}[x]\cdot(x-\mu)$ $\overset{?}{=}\hat{f}[\mu]-\tilde{p}[x]\cdot Z_{\mathcal{S}}[x]$ (16)
@@
TABLE XV: Enhanced Interactive zk-SNARK Protocol with Heterogeneous Data Sources and Batch Verification (${\tt S}_{\tt EV}$)
@@
    Prover $\Rightarrow$ Verifier: ${\mathcal{S}}_{X},\hat{k},\hat{s},\hat{s}_{1},\hat{s}_{2}$ computing is outsourced to prover:
@@
[…]
@@
TABLE XVI: Settings of vectors $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ in the constraint system for parametric bushfire insurance. $\mathbf{I}(i,j)$ represents a vector of length $(10n+(n+2)k)$ for which every entry are 0 except $i_{th},j_{th}$ entries which are 11.
@@
$\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i},i\in\{1,...,n\}$ 0 $\mathbf{I}(6n+(n+2)k+i,7n+(n+2)k+i)$ $\mathbf{-I}(i)$ $\mathbf{I}$
$\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i}-\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$ 0 $\mathbf{I}(6n+(n+2)k+i)-I(7n+(n+2)k+i,3n+i)$ $\mathbf{I}$ $\mathbf{-I}(i)$
$\mathbf{r}^{+}_{i}+\mathbf{s}^{+}_{i},i\in\{1,...,n\}$ 0 $\mathbf{I}(8n+(n+2)k+i,9n+(n+2)k+i)$ $\mathbf{-I}(n+i)$ $\mathbf{I}$
$\mathbf{r}^{+}_{i}-\mathbf{s}^{+}_{i}-\mathbf{\theta}^{+}_{i},i\in\{1,...,n\}$ 0 $\mathbf{I}(8n+(n+2)k+i)-I(9n+(n+2)k+i,4n+i)$ $\mathbf{I}$ $\mathbf{-I}(n+i)$
@@
$\sum^{k}_{j=1}e_{i,j},i\in\{1,...,n\}$ 0 $\mathbf{I}(6n+(i-1)k+1,...,6n+ik)$ $\mathbf{I}$ $\mathbf{-I}(2n+i)$
@@
Checking $\mathbf{i}_{i}$ in $\mathbf{a},\mathbf{b},i\in\{1,...,n\}$ 11 $\mathbf{I}(5n+i)$ $\mathbf{I}(5n+i)$ $\mathbf{I}$
@@
VIII-F Enhanced Sonic Protocol with Batch Verification
By replacing the original restricted KZG polynomial commitment scheme by the one with batch verification, we present an enhanced version of zk-SNARK protocol (${\tt S}_{\tt EV}$) in Table XV. To convert the interactive protocol ${\tt S}_{\tt EV}$ to be non-interactive, we can employ Fiat-Shamir heuristic to replace the verifier-supplied random challenges $(y,\beta,z)$ by hash values from the previous commitments.
VIII-G Implementation of Bushfire Insurance by Sonic Protocol
In this section, we describe how to encode the bushfire detection model in Sec.VI-A into a Sonic-compatible constraint system. We specify the detailed settings of Sonic-specific vectors $({\bf a},{\bf b},{\bf c})$ and $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$ in Tables XVI and XVII. We assume image size $n$ and $k$ is the length of bit indices.
We first construct the input vectors $({\bf a},{\bf b},{\bf c})$. Row $1$ to $n$ of these vectors are $\mathbf{n}^{-}$, $(\mathbf{r}^{-}+\mathbf{s}^{-})$ and $(\mathbf{r}^{-}-\mathbf{s}^{-}-\mathbf{\theta}^{-})$ respectively. Because ${\bf a}\cdot{\bf b}={\bf c}$, it verifies $(\mathbf{r}^{-}_{i}-\mathbf{s}^{-}_{i})=\mathbf{n}^{-}_{i}(\mathbf{r}^{-}_{i}+\mathbf{s}^{-}_{i})+\mathbf{\theta}^{-}_{i},i\in\{1,...,n\}$. Similarly, we construct row $n+1$ to $6n$ to verify the rest of the multiplication gates, which are shown in Table XVII. Row $5n+1$ to row $6n$ verifies $\bf i$ is binary and row $6n+1$ to $6n+(n+2)k$ verifies $\bf e,\theta_{d},G$ are a non-negative. Finally, row $6n+(n+2)k+1$ to the end is input from the data source.
Next we construct $({\mathbf{u}_{q},\mathbf{v}_{q},\mathbf{w}_{q}},k_{q})$. The first step is to verify entries in $({\bf a},{\bf b},{\bf c})$ are consistent with the input from data sources. For example, row $1$ t $n$ of $\bf b$ are $\mathbf{r}^{-}+\mathbf{s}^{-}$, while $\mathbf{r}^{-}$ and $\mathbf{s}^{-}$ are row 6n+(n+2)k+1,,8n+(n+2)k6n+(n+2)k+1,...,8n+(n+2)k of $\bf a$. Then the first constraint in Table XVI checks $\bf a_{6n+(n+2)k+i}+\bf a_{7n+(n+2)k+i}=\bf b_{i},i\in{1,...,n}$. The next step is to encode the sum gates in the circuits. For example, checking $\sum^{n}_{i=1}\mathbf{i}_{i}-G=\epsilon$, where $\bf i$ locates in row $5n+1$ to $6n$, and $G$ in row $6n+(n+1)k+1$ to $6n+(n+2)k$ in binary decomposition form, and $\epsilon$ in $k_{q}$ as a public input. The last part checks binary decompositions are correct.